OpenStudy (anonymous):

Find all solutions to the equation:: (sin x- cos x)^2=3

OpenStudy (reemii):

your first exercise told you that (sin x - cos x)^2 = 1 - sin(2x), so use that and transform the equation.

OpenStudy (anonymous):

So would we put 1-sin(2x) equal to 3? Because you can cross out the first part of the equation with the part you just examples

OpenStudy (reemii):

"So would we put 1-sin(2x) equal to 3" <- exactly, so after that you can use usual transformations to isolate "sin(2x)".

OpenStudy (anonymous):

So then you would have -sin(2x)=4 and then times by a negaitve giving you sin(2x)=-4. Then you can plug it into your calculator and figure out the answers?

OpenStudy (reemii):

in these type of equations, when only one trigonometric function remains, you usually want to isolate it, because something that looks like "sin(stuff) = 1" tells you something about "stuff".

OpenStudy (anonymous):

We should take the inverse of both sides and then 2x= sin^-1(4) right?

OpenStudy (anonymous):

Then you would divide by 2 and then i'm lost....

OpenStudy (reemii):

wait, 1-sin(2x) = 3 <=> -sin(2x) = 2 <=> sin(2x) = -2. Here, it looks like there's nothing to compute anymore: what you got is "sin(something) = -2", that is impossible.

OpenStudy (anonymous):

Ok, so then we will need to leave it that way.... A sin product can be negative in this case?

OpenStudy (anonymous):

actually it can be negative; but the thing is that sin value cannot exceed 1 and hence in any other quadrant, can also not be less than -1. so no solution.

OpenStudy (reemii):

what is "sin product" ? hmm, what I see is that "sin(whatever)" can never be equal to -2, it always some number between -1 and 1. Therefore you say: "no solution".

OpenStudy (reemii):

he just said it lol.

OpenStudy (anonymous):

Yea, great minds think alike. So for the next question it asks: find all solutions for (sin x- cos x)^2=1.....

OpenStudy (anonymous):

this one is quite simple. you can expand again and solve; come till where you're left only with 2 sin x cos x.

OpenStudy (anonymous):

That one wouldn't work either then correct? because -sin(2x) can't equal 0... because I put 1-sin(2x)=1 again like before and then subtraacted one

OpenStudy (anonymous):

LOL of course sin 2x can equal 0. bcoz 0 is sin 0 so sin 2x = sin 0 2x = 0.

OpenStudy (anonymous):

Ooook, but you just said that sin has to either equal -1 or 1 to be a solution... So 0 can work as well for these problems?>

OpenStudy (anonymous):

@HemerickLee he said BETWEEN -1 and 1.

OpenStudy (anonymous):

Ahhh, i'm sorry wrong reading.... Makes since now... So they can only equal -1,0,1 to be a solution available

OpenStudy (reemii):

when seeing "sin(2x)=0", all you need to thing of first is "can sin(soemthing) be equal to 0"? If the answer is yes (it is), then you conclude something about "something". "sin(something) = 0" means "something=0 or pi or 2pi ... or -pi or -2pi or ..". Correctly: \(\sin (2x) = 0 \iff 2x = k\pi,\) for \(k\in\mathbb Z (\text{integers}) \iff x = \frac{k\pi}{2}\) with \(k\) integer.

OpenStudy (anonymous):

@reemii so isn't it simply 0. ?

OpenStudy (anonymous):

So x= \[\frac{ \pi n }{ 2 }\]

OpenStudy (reemii):

there's more than just 0 as solution, since sin(0) = sin(2pi) = sin(4pi) etc..

OpenStudy (anonymous):

Like my answer says... It can be any thing with pi after it?

OpenStudy (reemii):

"x = n pi / 2" means, every x of the form \(n\pi/2\) (with \(n\) integer) is a solution.

OpenStudy (reemii):

what do you mean by "any thing with pi after it" ? (sounds dangerous..)

OpenStudy (anonymous):

Alright, Thank you guys! psssswwwwhhh glad to get that off of my chest...